# Fair Resource Allocation in Multi-Task Learning

### v1

@@ Abstract
By jointly learning multiple tasks, multi-task learning (MTL) can leverage the shared knowledge across tasks, resulting in improved data efficiency and generalization performance. However, a major challenge in MTL lies in the presence of conflicting gradients, which can hinder the fair optimization of some tasks and subsequently impede MTL’s ability to achieve better overall performance. Inspired by fair resource allocation in communication networks, we formulate the optimization of MTL as a utility maximization problem, where the loss decreases across tasks are maximized under different fairness measurements. To solve this problem, we propose FairGrad, a novel MTL optimization method. FairGrad not only enables flexible emphasis on certain tasks but also achieves a theoretical convergence guarantee. Extensive experiments demonstrate that our method can achieve state-of-the-art performance among gradient manipulation methods on a suite of multi-task benchmarks in supervised learning and reinforcement learning. Furthermore, we incorporate the idea of $\alpha$-fairness into loss functions of various MTL methods. Extensive empirical studies demonstrate that their performance can be significantly enhanced. Code is provided at https://github.com/OptMN-Lab/fairgrad.
@@ 1 Introduction
By aggregating labeled data for various tasks, multi-task learning (MTL) can not only capture the latent relationship across tasks but also reduce the computational overhead compared to training individual models for each task [1; 2; 3]. As a result, MTL has been successfully applied in various fields like natural language processing [4; 5; 6], computer vision [7; 8; 9], autonomous driving [10; 11; 12], and recommendation systems [13; 14; 15]. Research has shown that MTL is capable of learning robust representations, which in turn helps avoid overfitting certain individual tasks [16; 17; 18; 19], and hence often achieves better generalization than the single-task counterparts.
@@
MTL often solves the average loss across tasks in many real-world scenarios. However, it has been shown that there may exist conflicting gradients [20; 21; 22; 23] among tasks that exhibit different directions and magnitudes. If directly optimizing the average loss, the final update direction will often be dominated by the largest gradient, which can degrade the overall performance of MTL. To alleviate this negative impact, a series of gradient manipulation methods have been proposed to find a compromised direction [24; 25; 20; 21; 26; 27; 28]. In this paper, we view these methods from a novel fairness perspective. For example, MGDA [24] and its variants such as [27; 29; 21; 28] tends to strike a max-min fairness among tasks, where the least-fortune tasks (i.e., with the lowest progress) are the most important tasks. Nash-MTL [26] aims to achieve a proportional fairness among tasks by formulating the problem as a bargaining game, and aims to attain a balanced solution that is not dominated by any single large gradient.
@@
However, different applications may favor different types of task fairness, and there is currently no unified framework in MTL that allows for the incorporation of diverse fairness concepts beyond those previously mentioned. To fill this gap, we propose a novel fair MTL framework, as well as efficient algorithms with performance guarantee. Our specific contributions are summarized as below.
@@
  • We first draw an important connection between MTL and fair resource allocation in communication networks [30; 31; 32; 33; 34; 35; 36], where we think of the common search direction $d$ shared by all tasks as a resource to minimize their losses, and the service quality is measure by the loss decrease after performing a gradient descent along $d$. Inspired by this connection, we model MTL as a utility maximization problem, where each task is associated with $\alpha$-fair utility function and different $\alpha$ yields different ideas of fairness including max-min, proportional, minimum potential delay fairness, etc.
@@
  • We propose a novel algorithm named FairGrad to solve the $\alpha$-fair MTL utility maximization problem. FairGrad is easy to implement, allows for a flexible selection of $\alpha$, and guarantees a convergence to a Parato stationary point under mild assumptions.
@@
  • Finally, we incorporate our idea of $\alpha$-fairness into the loss functions of existing methods including linear scalarization, RLW, DWA, UW, MGDA, PCGrad and CAGrad, and demonstrate that it can significantly improve their overall performance.
@@
(b) FairGrad (proportional fair)
@@
Figure 1: An illustrative two-task example from [26] to show the convergence of FairGrad to Pareto front from different initialization points. The illustration showcases four fairness concepts (from left to right): simple average (i.e., linear scalarization (LS)), proportional fairness, minimum potential delay (MPD) fairness and max-min fairness. It can be seen that LS is inclined towards the task $2$ with a larger gradient. FairGrad with proportional fairness resembles Nash-MTL [26], and can find more balanced solutions along the Pareto front. MPD fairness aims to minimize the overall time for all tasks to converge, and shifts slightly more attention to some struggling tasks with smaller gradients. Max-min fairness emphasizes more on the less-fortune task with a smaller gradient magnitude. Also observe that our FairGrad ensures the convergence to the Pareto front from all different initialization points.
@@ 2 Related Work
Multi-Task Learning. MTL has drawn significant attention both in theory and practice. One class of studies is designing sophisticated model architectures. These studies can be mainly divided into two categories, hard parameter sharing where task-specific layers are built on a common feature space [37; 38], and soft parameter sharing which couples related parameters through certain constraints [39; 40]. Another line of research aims to capture the relationship among tasks to guide knowledge transfer effectively [41; 42]. Additionally, the magnitudes of losses for different tasks may vary, posing challenges to the optimization of MTL. A group of studies seeks to balance tasks through heuristic re-weighting rules such as task-dependent uncertainty [43], gradient magnitudes [25], and the rate of change of loss for each task [37].
@@
As MTL is one of the important applications of multi-objective optimization (MOO), several MOO-based gradient manipulation methods have been explored recently to address the challenge of conflicting gradients. [24] proposed MGDA, and show it guarantees the convergence to the Pareto front under certain assumptions. [23] casted MTL as a MOO problem and refined MGDA for optimization in the context of deep neural networks. [20] determined the update by projecting a task’s gradient onto the normal plane of other conflicting gradients. [21] limited the update to a neighborhood of the average gradient. [26] considered finding the update as a bargaining game across all tasks. [28] searched for the update with the largest worst-case loss improvement rate to ensure that all tasks are optimized with approximately similar progress.
@@
Theoretically, [44] showed that linear scalarization cannot fully explore the Pareto front compared with MGDA-variant methods. [45] proposed a correlation-reduced stochastic gradient manipulation method to address the non-convergence issue of MGDA, CAGrad and PCGrad in the stochastic setting. [29] introduced a stochastic variant of MGDA with guaranteed convergence. [27] proposed a simple and provable SGD-type method that benefits from direction-oriented improvements like [21]. [46] offered a framework for analyzing stochastic MOO algorithms, considering the trade-off among optimization, generalization, and conflict-avoidance.
@@
Fairness in Resource Allocation. Fair resource allocation has been studied for decades in wireless communication [47; 48; 49; 50; 51; 52], where limited resources such as power and communication bandwidth need to be fairly allocated to users of the networks. Various fairness criteria have been proposed to improve the service quality for all users without sacrificing the overall network throughput. For example, Jain’s fairness index [30] prefers all users to share the resources equally. Proportional fairness [31] distributes resources proportional to user demands or priorities. Max-min fairness [33] attempts to protect the user who receives the least amount of resources by providing them with the maximum possible allocation. The $\alpha$-fairness framework was proposed to unify multiple fairness criteria, where different choices of $\alpha$ lead to different ideas of fairness [32; 49]. Recent research has explored the application of fair resource allocation in federated learning [53; 54]. In this paper, we connect MTL with the problem of fair resource allocation, and further propose an $\alpha$-fair utility maximization problem as well as an efficient algorithm to solve it.
@@ 3.1 Multiple Objectives and Pareto Concepts
Given two points $\theta_{1},\theta_{2}\in\mathbb{R}^{m}$, we say that $\theta_{1}$ dominates $\theta_{2}$ if $l_{i}(\theta_{1})\geq l_{i}(\theta_{2})$ for all $i\in[K]$ and $L(\theta_{1})\neq L(\theta_{2})$. A point is considered Pareto optimal if it is not dominated by any other point. This means that no improvement can be made in any one objective without negatively affecting at least one other objective. The set of all Pareto optimal points form into the Pareto front. A point $\theta\in\mathbb{R}^{m}$ is Pareto stationary if $\min_{w\in\mathcal{W}}\|G(\theta)w\|=0$, where $G(\theta)=[g_{1}(\theta),\cdots,g_{K}(\theta)]\in\mathbb{R}^{m\times K}$ is the matrix with each column $g_{i}(\theta)$ denoting the gradient of $i$-th objective, and $\mathcal{W}$ is the probability simplex defined on $[K]$. Pareto stationarity is a necessary condition for Pareto optimality.
@@ 3.2 $\alpha$-Fair Resource Allocation
In the context of fair resource allocation in communication networks with $K$ users, the goal is to properly allocate resource (e.g., channel bandwidth, transmission rate) to maximize the total user utility (e.g., throughput) under the link capacity constraints. A generic overall objective for the network is given by
@@
Minimum Potential Delay Fairness. When $\alpha=2$, the utility function is $-\frac{1}{x_{i}}$ and hence the overall objective is to minimize $\sum_{i}\frac{1}{x_{i}}$. Since $x_{i}$ is the transmission rate in networks, $\frac{1}{x_{i}}$ can be reviewed as the delay of transferring a file with unit size, and hence this case is called as minimum potential delay fairness.
@@ 4 FairGrad: Fair Resource Allocation in MTL
Inspired by the fair resource allocation over networks in Section 3.2, we now provide an $\alpha$-fair framework for MTL. Let $d$ be the updating direction for all tasks within the ball $B_{\epsilon}$ centered at 00 with a radius of $\epsilon$, and $g_{i}$ be the gradient for task $i$. Then, based on the first-order Taylor approximation of $l_{i}(\theta)$, we have for a small stepsize $\tau$
$\frac{1}{\alpha}[l_{i}(\theta)-l_{i}(\theta-\alpha g_{i}^{\top}d)]\approx g_{i}^{\top}d,$
@@
where $\alpha\in[0,1)\cup(1,+\infty)$.
Note that our $\alpha$-fair framework takes the same spirit as Linear Scalarization (LS), Nash-MTL and MGDA when we take $\alpha\to 0,1,\infty$, respectively, and provides the coverage over other fairness ideas.
We next take the following steps to solve the problem in eq. 2. First note that the objective function is non-decreasing with respect to any feasible $d$. Thus, if $d$ lies in the interior of $B_{\epsilon}$, then there must exist a point along the same direction but on the boundary of $B_{\epsilon}$, which achieves a larger overall utility. Thus, it can be concluded that the optimal $d^{*}$ lies on the boundary and the gradient of the overall objective is aligned with $d^{*}$, i.e.,
@@
for some constant $c>0$. Following Nash-MTL [26], we take $c=1$ for simplicity, and assume that the gradients of tasks are linearly independent when not at a Pareto stationary point $\theta$ such that $d$ can be represented as a linear combination of task gradients: $d=\sum_{i}w_{i}g_{i}$, where $w:=(w_{1},...,w_{K})^{\top}\in\mathbb{R}_{+}^{K}$ denotes the weights. Then, we obtain from eq. 3 that $(g_{i}^{\top}d)^{-\alpha}=w_{i}$, which combined with $d=\sum_{i}w_{i}g_{i}$, implies that
@@
where $\alpha\neq 0$ and the power $-1/\alpha$ is applied elementwisely. It is evident that we have $w_{i}=1$ for all $i\in[K]$ when $\alpha=0$. Differently from Nash-MTL that approximates the solution using a sequence of convex optimization problems, we treat eq. 4 as a simple constrained nonlinear least square problem

### v2

@@ Abstract
By jointly learning multiple tasks, multi-task learning (MTL) can leverage the shared knowledge across tasks, resulting in improved data efficiency and generalization performance. However, a major challenge in MTL lies in the presence of conflicting gradients, which can hinder the fair optimization of some tasks and subsequently impede MTL’s ability to achieve better overall performance. Inspired by fair resource allocation in communication networks, we formulate the optimization of MTL as a utility maximization problem, where the loss decreases across tasks are maximized under different fairness measurements. To address the problem, we propose FairGrad, a novel optimization objective. FairGrad not only enables flexible emphasis on certain tasks but also achieves a theoretical convergence guarantee. Extensive experiments demonstrate that our method can achieve state-of-the-art performance among gradient manipulation methods on a suite of multi-task benchmarks in supervised learning and reinforcement learning. Furthermore, we incorporate the idea of $\alpha$-fairness into the loss functions of various MTL methods. Extensive empirical studies demonstrate that their performance can be significantly enhanced. Code is available at https://github.com/OptMN-Lab/fairgrad.
Multi-Task Learning, Fair Resource Allocation, Utility Maximization
@@ 1 Introduction
By aggregating labeled data for various tasks, multi-task learning (MTL) can not only capture the latent relationship across tasks but also reduce the computational overhead compared to training individual models for each task (Caruana, 1997; Evgeniou & Pontil, 2004; Thung & Wee, 2018). As a result, MTL has been successfully applied in various fields like natural language processing (Liu et al., 2016a; Zhang et al., 2023; Radford et al., 2019), computer vision (Zhang et al., 2014; Dai et al., 2016; Vandenhende et al., 2021), autonomous driving (Chen et al., 2018a; Ishihara et al., 2021; Yu et al., 2020a), and recommendation systems (Bansal et al., 2016; Li et al., 2020; Wang et al., 2020a). Research has shown that MTL is capable of learning robust representations, which in turn helps avoid overfitting certain individual tasks (Lounici et al., 2009; Zhang & Yang, 2021; Ruder, 2017; Liu et al., 2016b), and hence often achieves better generalization than the single-task counterparts.
@@
MTL often solves the average loss across tasks in many real-world scenarios. However, it has been shown that there may exist conflicting gradients (Yu et al., 2020b; Liu et al., 2021; Wang et al., 2020b; Sener & Koltun, 2018) among tasks that exhibit different directions and magnitudes. If directly optimizing the average loss, the final update direction will often be dominated by the largest gradient, which can degrade the overall performance of MTL. To alleviate this negative impact, a series of gradient manipulation methods have been proposed to find a compromised direction (Désidéri, 2012; Chen et al., 2018b; Yu et al., 2020b; Liu et al., 2021; Navon et al., 2022; Xiao et al., 2023; Liu et al., 2023). In this paper, we view these methods from a novel fairness perspective. For example, MGDA (Désidéri, 2012) and its variants such as (Xiao et al., 2023; Fernando et al., 2022; Liu et al., 2021, 2023) tends to strike a max-min fairness among tasks, where the least-fortune tasks (i.e., with the lowest progress) are the most important. Nash-MTL (Navon et al., 2022) aims to achieve proportional fairness among tasks by formulating the problem as a bargaining game, attaining a balanced solution that is not dominated by any single large gradient.
@@
However, different applications may favor different types of task fairness, and there is currently no unified framework in MTL that allows for the incorporation of diverse fairness concepts beyond those previously mentioned. To fill this gap, we propose a novel fair MTL framework, as well as efficient algorithms with performance guarantee. Our specific contributions are summarized below.
@@
  • We first draw an important connection between MTL and fair resource allocation in communication networks (Jain et al., 1984; Kelly, 1997; Mo & Walrand, 2000; Radunovic & Le Boudec, 2007; Srikant & Ying, 2013; Ju & Zhang, 2014; Liu & Xia, 2015), where we think of the common search direction $d$ shared by all tasks as a resource to minimize their losses, and the service quality is measure by the loss decrease after performing a gradient descent along $d$. Inspired by this connection, we model MTL as a utility maximization problem, where each task is associated with $\alpha$-fair utility function and different $\alpha$ yields different ideas of fairness including max-min, proportional, minimum potential delay fairness, etc.
@@
  • We propose a novel algorithm named FairGrad to solve the $\alpha$-fair MTL utility maximization problem. FairGrad is easy to implement, allows for a flexible selection of $\alpha$, and guarantees convergence to a Parato stationary point under mild assumptions.
@@
  • Finally, we incorporate our idea of $\alpha$-fairness into the loss functions of existing methods including Linear Scalarization, RLW, DWA, UW, MGDA, PCGrad, and CAGrad, and demonstrate that it can significantly improve their overall performance.
@@
(b) FairGrad (proportionally fair)
@@
Figure 1: An illustrative two-task example from (Navon et al., 2022) to show the convergence of FairGrad to Pareto front from different initialization points (black dots $\bullet$). The optimization trajectories are colored from orange to purple. The bold gray line represents the Pareto front. The illustration showcases four fairness concepts (from left to right): simple average (i.e., Linear Scalarization (LS)), proportional fairness, minimum potential delay (MPD) fairness, and max-min fairness. It can be seen that LS is inclined towards the task $2$ with a larger gradient. FairGrad with proportional fairness resembles Nash-MTL (Navon et al., 2022), and can find more balanced solutions along the Pareto front. MPD fairness aims to minimize the overall time for all tasks to converge, and shifts slightly more attention to some struggling tasks with smaller gradients. Max-min fairness emphasizes more on the less-fortune task with a smaller gradient magnitude. Also, observe that our FairGrad ensures the convergence to the Pareto front from all different initialization points.
@@ 2 Related Work
Multi-Task Learning. MTL has drawn significant attention both in theory and practice. One class of studies is designing sophisticated model architectures. These studies can be mainly divided into two categories, hard parameter sharing where task-specific layers are built on a common feature space (Liu et al., 2019; Kokkinos, 2017), and soft parameter sharing which couples related parameters through certain constraints (Ruder et al., 2019; Gao et al., 2020). Another line of research aims to capture the relationship among tasks to guide knowledge transfer effectively (Zhao et al., 2020; Ciliberto et al., 2017). Additionally, the magnitudes of losses for different tasks may vary, posing challenges to the optimization of MTL. A group of studies seeks to balance tasks through heuristic re-weighting rules such as task-dependent uncertainty (Kendall et al., 2018), gradient magnitudes (Chen et al., 2018b), and the rate of change of loss for each task (Liu et al., 2019).
@@
As MTL is one of the important applications of multi-objective optimization (MOO), several MOO-based gradient manipulation methods have been explored recently to address the challenge of conflicting gradients. (Désidéri, 2012) proposed MGDA, and show it guarantees the convergence to the Pareto front under certain assumptions. (Sener & Koltun, 2018) cast MTL as a MOO problem and refined MGDA for optimization in the context of deep neural networks. (Yu et al., 2020b) determined the update by projecting a task’s gradient onto the normal plane of other conflicting gradients. (Liu et al., 2021) limited the update to a neighborhood of the average gradient. (Navon et al., 2022) considered finding the update as a bargaining game across all tasks. (Liu et al., 2023) searched for the update with the largest worst-case loss improvement rate to ensure that all tasks are optimized with approximately similar progress.
@@
Theoretically, (Hu et al., 2023) showed that Linear Scalarization cannot fully explore the Pareto front compared with MGDA-variant methods. (Zhou et al., 2022) proposed a correlation-reduced stochastic gradient manipulation method to address the non-convergence issue of MGDA, CAGrad, and PCGrad in the stochastic setting. (Fernando et al., 2022) introduced a stochastic variant of MGDA with guaranteed convergence. (Xiao et al., 2023) proposed a simple and provable SGD-type method that benefits from direction-oriented improvements like (Liu et al., 2021). (Chen et al., 2023) offered a framework for analyzing stochastic MOO algorithms, considering the trade-off among optimization, generalization, and conflict-avoidance.
@@
Fairness in Resource Allocation. Fair resource allocation has been studied for decades in wireless communication (Nandagopal et al., 2000; Eryilmaz & Srikant, 2006; Lan et al., 2010; Huaizhou et al., 2013; Noor-A-Rahim et al., 2020; Xu et al., 2021), where limited resources such as power and communication bandwidth need to be fairly allocated to users of the networks. Various fairness criteria have been proposed to improve the service quality for all users without sacrificing the overall network throughput. For example, Jain’s fairness index (Jain et al., 1984) prefers all users to share the resources equally. Proportional fairness (Kelly, 1997) distributes resources proportional to user demands or priorities. Max-min fairness (Radunovic & Le Boudec, 2007) attempts to protect the user who receives the least amount of resources by providing them with the maximum possible allocation. The $\alpha$-fairness framework was proposed to unify multiple fairness criteria, where different choices of $\alpha$ lead to different ideas of fairness (Mo & Walrand, 2000; Lan et al., 2010). Recent research has explored the application of fair resource allocation in federated learning (Li et al., 2019; Zhang et al., 2022). In this paper, we connect MTL with fair resource allocation and further propose an $\alpha$-fair utility maximization problem as well as an efficient algorithm to solve it.
@@ 3.1 Multiple Objectives and Pareto Concepts
Given two points $\theta_{1},\theta_{2}\in\mathbb{R}^{m}$, we say that $\theta_{1}$ dominates $\theta_{2}$ if $l_{i}(\theta_{1})\leq l_{i}(\theta_{2})$ for all $i\in[K]$ and $L(\theta_{1})\neq L(\theta_{2})$. A point is considered Pareto optimal if it is not dominated by any other point. This means that no improvement can be made in any one objective without negatively affecting at least one other objective. The set of all Pareto optimal points form into the Pareto front. A point $\theta\in\mathbb{R}^{m}$ is Pareto stationary if $\min_{w\in\mathcal{W}}\|G(\theta)w\|=0$, where $G(\theta)=[g_{1}(\theta),\cdots,g_{K}(\theta)]\in\mathbb{R}^{m\times K}$ is the matrix with each column $g_{i}(\theta)$ denoting the gradient of $i$-th objective, and $\mathcal{W}$ is the probability simplex defined on $[K]$. Pareto stationarity is a necessary condition for Pareto optimality.
@@ 3.2 $\alpha$-Fair Resource Allocation
In the context of fair resource allocation in communication networks with $K$ users, the goal is to properly allocate resources (e.g., channel bandwidth, transmission rate) to maximize the total user utility (e.g., throughput) under the link capacity constraints. A generic overall objective for the network is given by
@@
Minimum Potential Delay Fairness. When $\alpha=2$, the utility function is $-\frac{1}{x_{i}}$ and hence the overall objective is to minimize $\sum_{i}\frac{1}{x_{i}}$. Since $x_{i}$ is the transmission rate in networks, $\frac{1}{x_{i}}$ can be reviewed as the delay of transferring a file with unit size, and hence this case is called minimum potential delay fairness.
@@ 4 FairGrad: Fair Resource Allocation in MTL
Inspired by the fair resource allocation over networks in Section 3.2, we now provide an $\alpha$-fair framework for MTL. Let $d$ be the updating direction for all tasks within the ball $B_{\epsilon}$ centered at 00 with a radius of $\epsilon$, and $g_{i}$ be the gradient for task $i$. Then, based on the first-order Taylor approximation of $l_{i}(\theta)$, we have for a small stepsize $\eta$
$\frac{1}{\eta}[l_{i}(\theta)-l_{i}(\theta-\eta d)]\approx g_{i}^{\top}d,$
@@
where $\alpha\in[0,1)\cup(1,+\infty)$. Note that our $\alpha$-fair framework takes the same spirit as Linear Scalarization (LS), Nash-MTL, and MGDA when we take $\alpha\to 0,1,\infty$, respectively, and provides the coverage over other fairness ideas.
4.1 Analogy
Utility. For each user $i$ in a communication network, it usually holds that the larger the allocated transmission rate $x_{i}$, the higher the user’s level of satisfaction. However, the total link capacity in a communication network is constrained. For each task $i$ in MTL scenarios, the larger the loss decreasing rate $g_{i}^{\top}d$, the more optimized the task becomes. As we consider the update direction $d$ within a ball $B_{\epsilon}$ centered at 0 with a radius of $\epsilon$, the feasible update progress for each task is also constrained. In a communication network, increasing the allocated transmission rate for one user may decrease the rate of other users. Similarly, conflicting gradients may occur in MTL.
Capacity Constraint. In the network resource allocation, the convex link capacity constraint refers to the constraints imposed on individual network links to ensure the packet transmission rate across each link does not exceed its capacity. It can be formulated as follows:
$\sum_{i\in L}x_{i}<C,$
where $x_{i}\geq 0$, $C$ represents the capacity of the network link, and $L$ denotes the users who transmit their packets on this link. In Equation 2, the loss decrease rate $g_{i}^{\top}d$ is modeled as a utility. The feasible update direction $d$ in the ball $B_{\epsilon}$. This is similar to the capacity constraint in network resource allocation because $d$ here cannot be arbitrarily large and then has capacity in MTL. In addition, the constraint on $d$ in our case is: $g_{i}^{\top}d\geq 0$ and $d\in B_{\epsilon}$ for all $i$, which turns out to be convex, as in network resource allocation where the capacity constraint is convex.
4.2 Method
We next take the following steps to solve the problem in Equation 2. First note that the objective function is non-decreasing with respect to any feasible $d$. Thus, if $d$ lies in the interior of $B_{\epsilon}$, then there must exist a point along the same direction but on the boundary of $B_{\epsilon}$, which achieves a larger overall utility. Thus, it can be concluded that the optimal $d^{*}$ lies on the boundary, and the gradient of the overall objective is aligned with $d^{*}$, i.e.,
@@ 4.2 Method
for some constant $c>0$. Following Nash-MTL (Navon et al., 2022), we take $c=1$ for simplicity, and assume that the gradients of tasks are linearly independent when not at a Pareto stationary point $\theta$ such that $d$ can be represented as a linear combination of task gradients: $d=\sum_{i}w_{i}g_{i}$, where $w:=(w_{1},...,w_{K})^{\top}\in\mathbb{R}_{+}^{K}$ denotes the weights. Then, we obtain from Equation 3 that $(g_{i}^{\top}d)^{-\alpha}=w_{i}$, which combined with $d=\sum_{i}w_{i}g_{i}$, implies that
@@
where $\alpha\neq 0$ and the power $-1/\alpha$ is applied elementwisely. It is evident that we have $w_{i}=1$ for all $i\in[K]$ when $\alpha=0$. Differently from Nash-MTL that approximates the solution using a sequence of convex optimization problems, we treat Equation 4 as a simple constrained nonlinear least square problem